OpenStudy (anonymous):

how to solve 2x-√15-4x=0

OpenStudy (anonymous):

if it is \[2x-\sqrt{15-4x}=0\] \[2x=\sqrt{15-4x}\]

OpenStudy (anonymous):

then do what after that

OpenStudy (anonymous):

then square both sides to get \[4x^2=15-4x\] \[4x^2+4x-15=0\] \[(2x-3)(2x+5)=0\] and so \[2x-3=0\] \[2x=3\] \[x=\frac{3}{2}\]

OpenStudy (anonymous):

or \[2x+5=0\] \[2x=-5\] \[x=-\frac{5}{2}\]

OpenStudy (anonymous):

but you have to check your answers because we squared as the second step. you need to plug in both of these numbers for x and see if they both work

OpenStudy (anonymous):

you can check for yourself but i think both work. is it clear how to check?

OpenStudy (anonymous):

hold on I'm checking it right now

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Yeah it works! thanks you're awesome!

OpenStudy (anonymous):

yw!